Nikki is PM contact, while Anne is on Holiday 15-30 July,
Anne Holiday 15-30 July, Nikki is PM cover
9/07/13 - Data refresh in TEST now happening on Wednesday /Thursday 10th - 11th July. A lot of problems on Friday with the Clone backup and on Monday in the time it was taking to do the refresh. - Snapshot backup happening Mon 8/07 night for work on Wednesday EST063 Space Audut upgrade project planned for Tuesday with refresh and deployment with external support backup - No progress with HMRC Final submission date - (which month), is currently Jule, which we have told HMRC we cannot meet
See journal note below for more detailed infomation.: https://www.projects.ed.ac.uk/project/hrs072/journal-17
Work outstanding
- HRTEST refresh still underway 10 July 2013
- HRLIVE snapshot completed (Wed am)
- Restore to TEST still with UNIX team (I130708-0625 ) WIP
- Config LIVE to TEST awaiting return of unidesk call - to do
- API rewrite work (Config)
- RTI Patches (2) done in DEV
- RTI patches needed in Test and Live (Config)
- RTI screen sort out for Payroll
- EAS Employee alignment submission - for Payroll / Finance for HMRC
- FPS Final Payment Submission - for Payroll / Finance for HMRC
- Additional configuration with Oracle consultant to complete
- HESA waiting to progress to TEST.
- HR waiting to train in TEST.
History of delay with Clone work planned for 4-5th July - now planned for 10-11th July
- Snapshot on Thursday, was found to miss a couple of files,
- Files restore took a long time
- Unix team delayed in starting the restore
- When started took a VERY long time
- On Monday Clone proces did not work, so now having to start again
- Redoing process from scratch Wed morning
- HRLIVE snapshot complete Wed AM (quickly)